# Information-Theoretic Thresholds for the Alignments of Partially Correlated Graphs

### v1

@@ Abstract
This paper studies the problem of recovering the hidden vertex correspondence between two correlated random graphs. We propose the partially correlated Erdős-Rényi graphs model, wherein a pair of induced subgraphs with a certain number are correlated. We investigate the information-theoretic thresholds for recovering the latent correlated subgraphs and the hidden vertex correspondence. We prove that there exists an optimal rate for partial recovery for the number of correlated nodes, above which one can correctly match a fraction of vertices and below which correctly matching any positive fraction is impossible, and we also derive an optimal rate for exact recovery. In the proof of possibility results, we propose correlated functional digraphs, which partition the edges of the intersection graph into two types of components, and bound the error probability by lower-order cumulant generating functions. The proof of impossibility results build upon the generalized Fano’s inequality and the recovery thresholds settled in correlated Erdős-Rényi graphs model.
Keywords— Graph alignments, information-theoretic thresholds, Erdős-Rényi random graphs, partial recovery, exact recovery
@@ 1 Introduction
Recently, there has been a surge in interest in the problems of detecting graph correlations and the alignments of two correlated graphs. These questions have emerged across various domains. For instance, in social networks, determining the similarity between friendship networks across different platforms has garnered attention [NS08, NS09].
In the realm of computer vision, where 3-D shapes are often represented as graphs with adjacency matrices, the identification of whether two graphs represent the same object holds significant importance in pattern recognition and image processing [BBM05, CSS06]. In computational biology, the representation of biological networks as graphs aids in understanding and quantifying their correlation [SXB08, VCL+11]. Furthermore, in natural language processing, the ontology alignment problem involves representing each sentence as a graph, with nodes denoting words. The task of determining whether a given sentence can be inferred from the text directly relates to graph matching problems [HNM05]. Numerous graph models exist, with the Erdős-Rényi random graph model being a prominent example, as proposed by [PA59] and [Gil59]:
@@
While there are inherent disparities between the Erdős-Rényi random graph model and networks derived from real-world scenarios, comprehensively understanding the Erdős-Rényi graphs remains profoundly significant. This understanding serves as a pivotal step in transitioning from solving detection and matching problems on Erdős-Rényi graphs to addressing challenges inherent in practical applications. The graph alignment problem entails identifying latent vertex correspondences between two graphs based on their structures. Following [PG11], for two random graphs $G_{1},G_{2}$ with vertex sets $V(G_{1}),V(G_{2})$ and edge sets $E(G_{1}),E(G_{2})$, a typical correlated graph model is correlated Erdős-Rényi random graph model:
@@
Let $\pi$ denote a latent bijective mapping from $V(G_{1})$ to $V(G_{2})$. We say a pair of graphs $(G_{1},G_{2})$ are correlated Erdős-Rényi graphs if both marginal distributions are $\mathcal{G}(n,p)$ and each pair of edges $(uv,\pi(u)\pi(v))$ for $u,v\in V(G_{1})$ follows the correlated bivariate Bernoulli distribution with correlation coefficient $\rho$.
Given observations on $G_{1}$ and $G_{2}$ under the correlated Erdős-Rényi graphs model, the goal is to recover the latent vertex mapping $\pi$. To quantify the performance of an estimator $\hat{\pi}$, we consider the following two recovery criterion:
@@
    |$|\{v\in V(G_{1}):\pi(v)=\hat{\pi}(v)\}|\geq\delta|V(G_{1})|.$ (1)
@@
    $\pi(v)=\hat{\pi}(v),\quad\forall~{}v\in V(G_{1}).$ (2)
The information-theoretic thresholds for partial and exact recoveries of $\pi$ between two correlated Erdős-Rényi graphs have been extensively studied in the recent literature.
  • Partial Recovery. [GML21] presented an impossibility result for partial recovery in the sparse regime characterized by constant average degree and correlation. [HM23] showed that $np(p\vee\rho)\gtrsim\log\left(1+\frac{\rho}{p}\right)\vee 1$ suffices for partial recovery, while $n\gtrsim d(p+\rho-p\rho\|p)\log n$ is necessary, where $d(p\|q)$ denotes the Kullback–Leibler (KL) divergence between Bernoulli distributions with mean $p$ and $q$, respectively. The recent work [WXY22] settled the sharp threshold for dense graphs with $\frac{p}{p\vee\rho}=n^{-o(1)}$ and the thresholds within a constant factor for sparse ones with $\frac{p}{p\vee\rho}=n^{-\Omega(1)}$. For the sparse case, [DD23b] proved a sharp threshold when $\frac{p}{p\vee\rho}=n^{-\alpha+o(1)}$ for $\alpha\in(0,1]$.
  • Exact Recovery. Based on the properties of the intersection graph under a permutation $\pi$, [CK16, CK17] showed that the Maximal Likelihood Estimator (MLE) achieves exact recovery and established an information-theoretical lower bound with a gap of $\omega(1)$. The results are sharpened by [WXY22] where the sharp threshold for exact recovery are derived.
While numerous studies have extensively investigated recovery procedures within the correlated Erdős-Rényi graphs model, it is however imperative to recognize that the signal present in many graph structures from realistic models is often inferior to that within the correlated Erdős-Rényi graph. This discrepancy emerges as many nodes in realistic graphs do not have corresponding nodes in the second correlated graphs. To offer a resolution to this concern, we propose the following model where on part of the nodes from two graphs are correlated.
@@
Let $S^{*}\subseteq V(G_{1})$ be a latent subset of vertices and $\pi^{*}$ be a latent injective mapping from $S^{*}$ to $V(G_{2})$. We say a pair of graphs $(G_{1},G_{2})$ are partially correlated Erdős-Rényi graphs if both marginal distributions are $\mathcal{G}(n,p)$ and each pair of edges $(uv,\pi^{*}(u)\pi^{*}(v))$ for $u,v\in S^{*}$ follows the correlated bivariate Bernoulli distribution with correlation coefficient $\rho$.
The case $S^{*}=V(G_{1})$ reduces to a pair of correlated Erdős-Rényi graphs in Definition 2. Under the model in Definition 3, given $S^{*}\subseteq V(G_{1})$ and the range of $\pi^{*}$ denoted by $T^{*}\subseteq V(G_{2})$, the induced subgraphs $G_{1}[S^{*}]$ and $G_{2}[T^{*}]$ are correlated Erdős-Rényi graphs on $m$ vertices. Therefore, the model can be equivalently constructed by planting correlated Erdős-Rényi graphs over a pair of independent Erdős-Rényi graphs.
In this paper, we investigate the information-theoretic thresholds for recovering the correlated nodes $S^{*}$ and the mapping $\pi^{*}$. For notational simplicity, we also refer to the problem as recovering $\pi^{*}$ while keeping $S^{*}$ implicit as the domain of $\pi^{*}$. The success criterion is similar to (1) and (2), where $V(G_{1})$ shall be replaced by $S^{*}$. However, due to the potential inconsistency between the domain of $\pi^{*}$ and the estimator $\hat{\pi}:\hat{S}\mapsto V(G_{2})$, we define their overlap by:
@@
With the notion of overlap, the success criterion is given by
@@ 1.1 Main Results
In this subsection, we present the main results of the paper. We first introduce some notations for the presentation of main theorems. Throughout the paper, we assume $0<\rho\leq 1$, $0<p\leq\frac{1}{2}$, and the cardinality $|S^{*}|=m$ is known. We further assume $p\geq\frac{1}{n}$ since otherwise partial recovery is impossible by [WXY22]. For a pair of Bernoulli random variables with means $p_{1},p_{2}$ and correlation $\rho$, their bivariate distribution is denoted as $\mathrm{Bern}(p_{1},p_{2},\rho)$. In our model, a pair of correlated edges $(e,\pi^{*}(e))\sim\mathrm{Bern}(p,p,\rho)$. Define $p_{ij}\triangleq\mathbb{P}\left[e=i,\pi^{*}(e)=j\right]$ for $i,j\in\{0,1\}$. Then
$p_{11}=p^{2}+\rho p(1-p),\quad p_{10}=p_{01}=(1-\rho)p(1-p),\quad p_{00}=(1-p)^{2}+\rho p(1-p).$
For a pair $(e,\pi^{*}(e))$ both edges present with probability $p_{11}$, while for $\pi(e)\neq\pi^{*}(e)$ both $e$ and $\pi(e)$ present with probability $p^{2}$. The relative signal strength present in correlated edges is denoted by $\gamma\triangleq\frac{p_{11}}{p^{2}}-1=\frac{\rho(1-p)}{p}$. It turns out that such reparametrization of the correlation coefficient is crucial in determining the fundamental limits of the graph alignment problem.
Let $\mathcal{S}_{n,m}$ denote the set of injective mappings $\pi:S\subseteq V(G_{1})\mapsto V(G_{2})$ with $|S|=m$. Our goal is to identify the minimum number of correlated nodes $m$ such that recovery of $\pi^{*}$ is possible. For the possibility results, we consider the estimator defined as
$\hat{\pi}=\mathop{\text{argmax}}\limits_{\pi\in\mathcal{S}_{n,m}}\sum_{u\neq v}\mathbbm{1}_{uv\in E(G_{1})}\mathbbm{1}_{\pi(u)\pi(v)\in E(G_{2})}.$ (4)
Next, we introduce our main theorems. Define $\phi(\gamma)\triangleq(1+\gamma)\log(1+\gamma)-\gamma$.
Theorem 1 (Partial recovery).
For any constant $\delta\in(0,1)$, there exists constant $c_{1}(\delta)$ such that, when $m\geq\frac{c_{1}(\delta)\log n}{p^{2}\phi(\gamma)}$, for any $\pi^{*}\in\mathcal{S}_{n,m}$, the estimator in (4) satisfies
$\mathbb{P}\left[\mathrm{overlap}(\pi^{*},\hat{\pi})\geq\delta\right]=1-o(1).$
Furthermore, for any $c\in(0,1)$, there exists $c_{2}(c,\delta)$ such that, when $m\leq\frac{c_{2}(c,\delta)\log n}{p^{2}\phi(\gamma)}$, for any estimator $\hat{\pi}$,
@@
The possibility result is presented in the minimax sense, while the impossibility result is under a Bayesian model. Hence, the threshold holds for both minimax and Bayesian risks. Theorem 1 implies, for the purpose of partial recovery, the threshold for the number of correlated nodes $m$ is of the order $\frac{\log n}{p^{2}\phi(\gamma)}$, beyond which partial recovery is possible and below which partial recovery is impossible. The dependency on the ambient graph order is only logarithmic, while the scale in terms of $p$ and $\rho$ is characterized by $\frac{1}{p^{2}\phi(\gamma)}$.
Theorem 2 (Exact recovery).
When $m\geq C\left(\frac{\log n}{p^{2}\phi(\gamma)}\vee\frac{\log(1/(p^{2}\gamma))}{p^{2}\gamma}\right)$, where $C$ is a universal constant, for any $\pi^{*}\in\mathcal{S}_{n,m}$, the estimator in (4) satisfies
$\mathbb{P}\left[\mathrm{overlap}(\pi^{*},\hat{\pi})=1\right]=1-o(1).$
Furthermore, for any $c\in(0,1)$, there exists a constant $c_{3}$ only depending on $c$ such that, when $m\leq c_{3}\left(\frac{\log n}{p^{2}\phi(\gamma)}\vee\frac{\log(1/(p^{2}\gamma))}{p^{2}\gamma}\right)$, for any estimator $\hat{\pi}$
@@
Theorem 2 implies, for the purpose of exact recovery, the threshold for the number of correlated nodes $m$ is of the order $\frac{\log n}{p^{2}\phi(\gamma)}\vee\frac{\log(1/(p^{2}\gamma))}{p^{2}\gamma}$. Under the weak signal regime $\gamma=O(1)$, we obtain the same rate as for partial recovery described in Theorem 1. Although the $\log n$ scaling has been observed in many other problems on random graphs, under the strong signal regime $\gamma=\omega(1)$, Theorem 2 highlights a transition from $\frac{\log n}{p^{2}\phi(\gamma)}$ to $\frac{\log(1/(p^{2}\gamma))}{p^{2}\gamma}$ if $\log^{2}\frac{1}{p}-\log^{2}\frac{1}{\rho}\gtrsim\log n$. In the latter regime, the difficulty is essentially the recovery of mapping given the sets of correlated nodes $(S^{*},T^{*})$. See more discussions in Section 4.
In comparison to prior work, our results of partial recovery in Theorem 1 match the thresholds established in [WXY22] up to a constant factor in both dense and sparse regimes for the special case $S^{*}=V(G_{1})$. Furthermore, the threshold $\frac{\log(1/(p^{2}\gamma))}{p^{2}\gamma}$ for exact recovery is derived from addressing the alignment problem for the subgraphs with the additional information on the domain and range of $\pi^{*}$, which applies the result in [WXY22].

### v2

@@ Abstract
This paper studies the problem of recovering hidden vertex correspondences between two correlated random graphs. We introduce the partially correlated Erdős-Rényi model and the partially correlated Gaussian Wigner model, where a pair of induced subgraphs is correlated. We investigate the information-theoretic thresholds for recovering these latent correlated subgraphs and their hidden vertex correspondences. For the partially correlated Erdős-Rényi model, we establish the optimal rate for partial recovery: above this threshold, a positive fraction of vertices can be correctly matched, while below it, matching any positive fraction is impossible. We also determine the optimal rate for exact recovery. In the partially correlated Gaussian Wigner model, the optimal rates for partial and exact recovery coincide. To prove the achievability results, we introduce correlated functional digraphs to partition the edges and bound error probabilities using lower-order cumulant generating functions. Our impossibility results rely on a generalized Fano’s inequality and the recovery thresholds for correlated Erdős-Rényi graphs.
Keywords— Graph alignments, information-theoretic thresholds, Erdős-Rényi model, Gaussian Wigner model, partial recovery, exact recovery
@@ 1 Introduction
Recently, there has been a surge of interest in the problems of detecting graph correlations and recovering the alignments of two correlated graphs. These questions have emerged across various domains. For instance, in social networks, determining the similarity between friendship networks across different platforms has garnered attention [NS08, NS09]. In the realm of computer vision, the identification of whether two graphs represent the same object holds significant importance in pattern recognition and image processing [BBM05, CSS06]. In computational biology, the representation of biological networks as graphs aids in understanding and quantifying their correlation [SXB08, VCL+15]. Furthermore, in natural language processing, the task of determining whether a given sentence can be inferred from the text directly relates to graph matching problems [HNM05].
Numerous graph models exist, with the Erdős-Rényi random graph model being a prominent example, as proposed by [ER59] and [Gil59]:
@@
While there are inherent disparities between the Erdős-Rényi graph and networks derived from real-world scenarios, comprehensively understanding the Erdős-Rényi graphs remains profoundly significant. This understanding serves as a pivotal step in transitioning from solving detection and matching problems on Erdős-Rényi graphs to addressing challenges inherent in practical applications. The graph alignment problem entails identifying latent vertex correspondences between two graphs based on their structures. Following [PG11], for two random graphs $G_{1},G_{2}$, a common graph model is the correlated Erdős-Rényi model. For a weighted graph $G$ with vertex set $V(G)$ and edge set $E(G)$, the weight associated with each edge $uv$ is denoted as $\beta_{uv}(G)$ for any $u,v\in V(G)$. For an unweighted graph $G$, we define $\beta_{uv}(G)={\mathbf{1}_{\left\{{uv\in E(G)}\right\}}}$.
@@
Let $\pi$ denote a latent bijective mapping from $V(G_{1})$ to $V(G_{2})$. We say a pair of graphs $(G_{1},G_{2})$ is correlated Erdős-Rényi graphs if both marginal distributions are $\mathcal{G}(n,p)$ and each pair of edges $(\beta_{uv}(G_{1}),\beta_{\pi(u)\pi(v)}(G_{2}))$ for $u,v\in V(G_{1})$ follows the correlated bivariate Bernoulli distribution with correlation coefficient $\rho$.
We note that the edges in the Erdős-Rényi model are binary 0-1 random variables, where $\beta_{uv}(G_{1}),\beta_{\pi(u)\pi(v)}(G_{2})\in\left\{0,1\right\}$ for any $u,v\in V(G_{1})$. The Bernoulli-based Erdős-Rényi model, while useful for modeling binary relationships, can be limited when we aim to represent weighted edges or capture more complex dependencies between nodes. Another important model is the correlated Gaussian Wigner model proposed in [DMWX21] as a prototypical model for random graphs, where the edges follow Gaussian distributions.
Definition 3 (Correlated Gaussian Wigner model).
Let $\pi$ denote a latent bijective mapping from $V(G_{1})$ to $V(G_{2})$. We say a pair of graphs $(G_{1},G_{2})$ follows correlated Gaussian Wigner model if each pair of weighted edges $(\beta_{uv}(G_{1}),\beta_{\pi(u)\pi(v)}(G_{2}))$ follows bivariate normal distribution $\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\right)$ for any vertices $u,v\in V(G_{1})$.
Given observations on $G_{1}$ and $G_{2}$ under the correlated Erdős-Rényi graphs model or correlated Gaussian Wigner model, the goal is to recover the latent vertex mapping $\pi$. Let $\mathrm{Supp}(\pi)$ denote the support set of $\pi$. To quantify the performance of an estimator $\hat{\pi}$, we consider the following two recovery criteria:
@@
    |$|\{v\in\mathrm{Supp}(\pi):\pi(v)=\hat{\pi}(v)\}|\geq\delta|\mathrm{Supp}(\pi)|.$ (1)
@@
    $\pi(v)=\hat{\pi}(v),\quad\forall~{}v\in\mathrm{Supp}(\pi).$ (2)
The information-theoretic thresholds for partial and exact recoveries of $\pi$ under correlated Erdős-Rényi model and correlated Gaussian Wigner model have been extensively studied in the recent literature.
  • Erdős-Rényi model, partial recovery. In the sparse regime where $np$ and $\rho$ are constant, partial recovery is impossible [GML21]. It is shown in [HM23] that $np(p\vee\rho)\gtrsim\log\left(1+\frac{\rho}{p}\right)\vee 1$ suffices for partial recovery, while $n\gtrsim d(p+\rho-p\rho\|p)\log n$ is necessary, where $d(p\|q)$ denotes the Kullback–Leibler (KL) divergence between Bernoulli distributions with mean $p$ and $q$, respectively. The recent work [WXY22] settled the sharp threshold for dense graphs with $\frac{p}{p\vee\rho}=n^{-o(1)}$ and the thresholds within a constant factor for sparse ones with $\frac{p}{p\vee\rho}=n^{-\Omega(1)}$. For the sparse case, a sharp threshold has been proven when $\frac{p}{p\vee\rho}=n^{-\alpha+o(1)}$ for $\alpha\in(0,1]$ in [DD23b].
  • Erdős-Rényi model, exact recovery. Based on the properties of the intersection graph under a permutation $\pi$, it is shown in [CK16, CK17] that the Maximal Likelihood Estimator (MLE) achieves exact recovery and establishes an information-theoretical lower bound with a gap of $\omega(1)$. The results are sharpened by [WXY22] where the sharp threshold for exact recovery are derived.
  • Gaussian Wigner model. It is shown in [Gan22] that if $n\rho^{2}\geq(4+\epsilon)\log n$ for any constant $\epsilon>0$, then the MLE achieves exact recovery; if instead $n\rho^{2}\leq(4-\epsilon)\log n$, then exact recovery is impossible. The results are strengthened by [WXY22] by showing that even partial recovery is impossible under the same condition.
While numerous studies have extensively investigated recovery procedures in correlated Erdős-Rényi and correlated Gaussian Wigner models, it is however imperative to recognize that, in real-world applications, many nodes in one graph may not have corresponding counterparts in the other graph, leading to incomplete or misaligned structural information. To offer a resolution to this concern, we propose the following models where only a subset of the nodes between the two graphs are correlated.
@@
Let $S^{*}\subseteq V(G_{1})$ be a latent subset of vertices and $\pi^{*}:S^{*}\mapsto V(G_{2})$ be a latent injective mapping. We say a pair of graphs $(G_{1},G_{2})$ is partially correlated Erdős-Rényi graphs if both marginal distributions are $\mathcal{G}(n,p)$ and each pair of weighted edges $(\beta_{uv}(G_{1}),\beta_{\pi^{*}(u)\pi^{*}(v)}(G_{2}))$ for $u,v\in S^{*}$ follows the correlated bivariate Bernoulli distribution with correlation coefficient $\rho$.
Definition 5 (Partially correlated Gaussian Wigner model).
Let $S^{*}\subseteq V(G_{1})$ be a latent subset of vertices and $\pi^{*}:S^{*}\mapsto V(G_{2})$ be a latent injective mapping. We say a pair of graphs $(G_{1},G_{2})$ follows partially correlated Gaussian Wigner model if the marginal distribution of each edge in both graphs is standard normal, and for $u,v\in S^{*}$, the pair $(\beta_{uv}(G_{1}),\beta_{\pi^{*}(u)\pi^{*}(v)}(G_{2}))$ follows bivariate normal distribution with correlation coefficient $\rho$.
Let $G[S]$ denote the induced subgraph of $G$ with vertex set $S\subseteq V(G_{1})$. For the partially models in Definitions 4 and 5, given $S^{*}\subseteq V(G_{1})$ and the range of $\pi^{*}$ denoted by $T^{*}=\pi^{*}(S^{*})\subseteq V(G_{2})$, the induced subgraphs $G_{1}[S^{*}]$ and $G_{2}[T^{*}]$ follow correlated Erdős-Rényi model and correlated Gaussian Wigner model on $m$ vertices, respectively. Specifically, the case $S^{*}=V(G_{1})$ reduces to correlated Erdős-Rényi model and correlated Gaussian Wigner model in Definitions 2 and 3.
In this paper, we investigate the information-theoretic thresholds for recovering the set of correlated nodes $S^{*}$ and the mapping $\pi^{*}$. For notational simplicity, we also refer to the problem as recovering $\pi^{*}$ while keeping $S^{*}$ implicit as the domain of $\pi^{*}$. The success criteria are given by (1) and (2), where $\mathrm{Supp}(\pi)=S^{*}$. Owing to the potential inconsistency between the domain of $\pi^{*}$ and that of the estimator $\hat{\pi}:\hat{S}\mapsto V(G_{2})$, we define their overlap by
@@
With the notion of overlap, the success criteria are equivalent to
@@ 1.1 Main Results
In this subsection, we present the main results of the paper. We first introduce some notations for the presentation of the main theorems. Throughout the paper, we assume that $10<\rho<$ 1, $0<p\leq\frac{1}{2}$, and the cardinality $|S^{*}|=m$ is known. We denote the bivariate distribution of a pair of Bernoulli random variables with means $p_{1},p_{2}$, and correlation coefficient $\rho$ as $\mathrm{Bern}(p_{1},p_{2},\rho)$. Specifically, for $\mathrm{Bern}(p,p,\rho)$, we denote the following probability mass function:
$p_{11}\triangleq p^{2}+\rho p(1-p),\quad p_{10}=p_{01}\triangleq(1-\rho)p(1-p),\quad p_{00}\triangleq(1-p)^{2}+\rho p(1-p).$ (4)
  • In the Erdős-Rényi model, a pair of correlated edges
    $\left(\beta_{uv}(G_{1}),\beta_{\pi^{*}(u)\pi^{*}(v)}(G_{2})\right)\sim\mathrm{Bern}(p,p,\rho).$
    Specifically, two correlated edges are both present with probability $p_{11}$, whereas two independent edges are both present with probability $p^{2}$. The relative signal strength is quantified by $\gamma\triangleq\frac{p_{11}}{p^{2}}-1=\frac{\rho(1-p)}{p}$. This reparametrization of the correlation coefficient is crucial in determining the fundamental limits of the graph alignment problem.
  • In the Gaussian Wigner model, a pair of correlated edges consists of two standard Gaussian random variables with correlation coefficient $\rho\in(0,1)$:
    $(\beta_{uv}(G_{1}),\beta_{\pi^{*}(u)\pi^{*}(v)}(G_{2}))\sim\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\right).$
    Here, the relative signal strength is directly characterized by the correlation coefficient $\rho$.
In the Erdős-Rényi model, we assume $p\geq\frac{1}{n}$, as partial recovery is otherwise impossible [WXY22]. Define
$\phi(\gamma)\triangleq(1+\gamma)\log(1+\gamma)-\gamma,\quad\gamma\triangleq\frac{\rho(1-p)}{p},$ (5)
and let $\mathcal{S}_{n,m}$ denote the set of injective mappings $\pi:S\subseteq V(G_{1})\mapsto V(G_{2})$ with $|S|=m$. Our goal is to determine the minimum number of correlated nodes $m$ required for successful recovery of $\pi^{*}$. Next, we introduce our main theorems.
Theorem 1 (Erdős-Rényi model, partial recovery).
There exists an estimator $\hat{\pi}$ such that, for any constant $\delta\in(0,1)$ and $\pi^{*}\in\mathcal{S}_{n,m}$,
$\mathbb{P}\left[\mathrm{overlap}(\pi^{*},\hat{\pi})\geq\delta\right]=1-o(1),$
when $m\geq\frac{c_{1}(\delta)\log n}{p^{2}\phi(\gamma)}$, where $c_{1}(\delta)$ is a constant depending on $\delta$.
Conversely, for any constant $c,\delta\in(0,1)$, there exists a constant $c_{2}(c,\delta)$ such that, when $m\leq\frac{c_{2}(c,\delta)\log n}{p^{2}\phi(\gamma)}$, for any estimator $\hat{\pi}$,
@@
The possibility result is established in the minimax sense, while the impossibility result is under a Bayesian framework. Hence, the threshold applies to both minimax and Bayesian risks. Theorem 1 implies that for partial recovery, the threshold for the number of correlated nodes $m$ is of the order $\frac{\log n}{p^{2}\phi(\gamma)}$, beyond which partial recovery is possible and below which partial recovery is impossible. The dependency on the ambient graph order $n$ is logarithmic, while the scaling with respect to $p$ and $\rho$ is characterized by $\frac{1}{p^{2}\phi(\gamma)}$.
Theorem 2 (Erdős-Rényi model, exact recovery).
There exists an estimator $\hat{\pi}$ such that, for any $\pi^{*}\in\mathcal{S}_{n,m}$,
$\mathbb{P}\left[\mathrm{overlap}(\pi^{*},\hat{\pi})=1\right]=1-o(1),$
when $m\geq C_{1}\left(\frac{\log n}{p^{2}\phi(\gamma)}\vee\frac{\log(1/(p^{2}\gamma))}{p^{2}\gamma}\right)$, where $C_{1}$ is a universal constant.
Conversely, for any $c\in(0,1)$, there exists a constant $c_{3}$ only depending on $c$ such that, when $m\leq c_{3}\left(\frac{\log n}{p^{2}\phi(\gamma)}\vee\frac{\log(1/(p^{2}\gamma))}{p^{2}\gamma}\right)$, for any estimator $\hat{\pi}$
@@
For exact recovery, Theorem 2 establishes the threshold for $m$ of the order $\frac{\log n}{p^{2}\phi(\gamma)}\vee\frac{\log(1/(p^{2}\gamma))}{p^{2}\gamma}$. Under the weak signal regime where $\gamma=O(1)$, this rate coincides with that for partial recovery in Theorem 1. While the logarithmic scaling in $n$ is common to many other problems on random graphs, under the strong signal regime where $\gamma=\omega(1)$, Theorem 2 reveals a transition from $\frac{\log n}{p^{2}\phi(\gamma)}$ to $\frac{\log(1/(p^{2}\gamma))}{p^{2}\gamma}$ when $\log^{2}\frac{1}{p}-\log^{2}\frac{1}{\rho}\gtrsim\log n$. In this regime, the challenge is essentially the oracle recovery of mapping given the sets of correlated nodes $(S^{*},T^{*})$. See more discussions in Section 4.
Theorem 3 (Gaussian Wigner model).
For any $\rho\in(0,1)$, there exists an estimator $\hat{\pi}$ such that, for any $\pi^{*}\in\mathcal{S}_{n,m}$,
$\mathbb{P}\left[\mathrm{overlap}(\pi^{*},\hat{\pi})=1\right]=1-o(1),$
when $m\geq C_{2}\left(\frac{\log n}{\log\left(1/(1-\rho^{2})\right)}\vee 1\right)$, where $C_{2}$ is a universal constant.
Conversely, for any constant $c,\delta\in(0,1)$, there exists a constant $c_{4}(c,\delta)$ such that, when $m\leq c_{4}(c,\delta)\left(\frac{\log n}{\log\left(1/(1-\rho^{2})\right)}\vee 1\right)$, for any estimator $\hat{\pi}$,
where $\pi^{*}$ is uniformly distributed over $\mathcal{S}_{n,m}$.
Theorem 3 implies that in the Gaussian Wigner model, the thresholds for the number of correlated nodes $m$ are of the order $\frac{\log n}{\log\left(1/(1-\rho^{2})\right)}\vee 1$, beyond which exact recovery is possible and below which partial recovery is impossible. There is no gap between partial and exact recovery. Indeed, the estimation of $\pi^{*}$ can be divided into two steps: identifying the correlated nodes $(S^{*},T^{*})$ and recovering the mapping given $(S^{*},T^{*})$. Since the relative signal strength is characterized by the bounded correlation coefficient $\rho$, identifying the correlated nodes is always the major challenge, with no gap existing under the two recovery regimes.
Remark 1.
The optimal rates for partial and exact recoveries under both correlated Erdős-Rényi model and correlated Gaussian Wigner model are derived in [WXY22]. Our results applied with $S^{*}=V(G_{1})$ match the thresholds established in their work up to a constant factor. Furthermore, the lower bound $\frac{\log(1/(p^{2}\gamma))}{p^{2}\gamma}$ for exact recovery is derived from addressing the alignment problem for the subgraphs with the additional information on the domain and range of $\pi^{*}$, which applies the result in their work.